Solve the following equations using Wronskian’s method y ՜ ՜ -3y ՜ +2y = e^2t /(e^t +1).
1
Expert's answer
2020-10-13T17:04:44-0400

"\displaystyle y'' - 3y' + 2y = \frac{e^{2t}}{e^{t} + 1} \\

\textsf{The solution to the above equation is}\\
y = y_c + y_p \\

\textsf{where}\, y_c \, \textsf{is the complementary factor and}\\
 y_p\, \textsf{ is the particular integral}\\

\textsf{The auxiliary equation is} \,m^2 - 3m + 2 = 0 \\

\begin{aligned}
(m - 1)(m - 2) &= 0 \\
\therefore m &= 1, 2
\end{aligned}\\
\therefore y_c = C_1e^t + C_2e^{2t} \\

\textsf{The Wronskian of the two solutions is}\\

W(t) = 
\begin{vmatrix}
e^t & e^{2t} \\
\frac{\mathrm{d}}{\mathrm{d}t}(e^t) & \frac{\mathrm{d}}{\mathrm{d}t}(e^{2t})
\end{vmatrix} = \begin{vmatrix}
e^t & e^{2t} \\
e^t & 2e^{2t}
\end{vmatrix} = 2e^{3t} - e^{3t} = e^{3t} \\

\textsf{Our particular solution will be given by}\\

y_p = V_1(t)e^{t} +V_2(t)e^{2t}\\

\textsf{Where}\, V_1(t) = -\int \frac{r(t)e^{2t}}{W(t)}\, \mathrm{d}t \\

\begin{aligned}
V_1(t) &= -\int \frac{e^{2t}}{e^t + 1} \cdot \frac{e^{2t}}{e^{3t}}\,\mathrm{d}t \\
&= -\int \frac{e^{t}}{e^t + 1} \,\mathrm{d}t = -\ln(e^t + 1) + C
\end{aligned}\\

\begin{aligned}
V_2(t) = \int \frac{r(t)e^{t}}{W(t)}\, \mathrm{d}t \\
&= \int \frac{e^{2t}}{e^t + 1} \cdot \frac{e^{t}}{e^{3t}}\,\mathrm{d}t\\
&= \int \frac{1}{e^t + 1} \,\mathrm{d}t = \ln\left(\frac{e^t}{e^t + 1}\right) + C
\end{aligned} \\

\textsf{The constant terms of the integration can be}\\\textsf{ignored since we are trying to find a non-constant}\\\textsf{solution to the differential equation}\\
\begin{aligned}
\therefore y &= y_c + y_p 
\\&= C_1e^t + C_2e^{2t} -\ln(e^t + 1)e^{t} +\ln\left(\frac{e^t}{e^t + 1}\right)e^{2t}
\end{aligned}"